OpenStudy (anonymous):

The speed of Train A is 18 mph slower than the speed of train B. Train A travels 210 miles in the same time it takes B to travel 300 miles. Find the speed of each train.

OpenStudy (anonymous):

Let x and y be the speeds of A and B. (x,y) is the solution to the simultaneous equations (1) y - x = 18 (2) 210/x = 300/y From equation (2): 300/y = 210/x = (300 - 210)/(y - x) Plugging in the value of y - x given in (1) 300/y = 210/x = 90/18 300/y = 210/x = 5 Hence Speed of train A: x = 210/5 = 42 mph Speed of train B: y = 300/5 = 60 mph